OpenStudy (anonymous):

Can anyone please help me with the following: Determine the fourth roots of -16 in the form x+iy Thank you

OpenStudy (anonymous):

-16 = 16* (-1) You want the 4th roots (as in 4th roots of unity) or the fourth root (1+i)/sqrt2

OpenStudy (anonymous):

The fourth roots of -16 in the form x+iy, please.

OpenStudy (anonymous):

Put -1 = 1(cos pi + i sin pi) in polar and use de Moivre

OpenStudy (anonymous):

U have -16 = 16 * (-1) so 4th root of 16 is 2 and then use de Moivre on the -1

OpenStudy (anonymous):

So, consider some polar number \(n\) with coordinates \(r, \theta\), we then have, by Euler: \[ n=r(\cos\theta+i\sin\theta)=r\,\text{cis}\,\theta=re^{i\theta} \]I'll be using the latter formula as opposed to the other two, simply because it's easier to type. So, by Euler's formula (or an extension to de Moivre's), \(n^\frac{1}{l}\) is: \[ \sqrt[l]{n}=\sqrt[l]{re^{i\theta}}=\left(\sqrt[l]{r}\right)\left(e^{i\left(\frac{\theta+2k\pi}{l}\right)}\right) \]For some integer \(k\). So, we apply this, to find: \[ n=16e^{i\pi} \]So: \[ \sqrt[4]n=2e^{i\frac{\pi}{4}}=2\left(\cos\left(\frac{\pi}{4}\right)+i\sin\left(\frac{\pi}{4}\right)\right)=\sqrt{2}+i\sqrt{2}\\ \sqrt[4]n=2e^{i\frac{\pi+2\pi}{4}}=2\left(\cos\left(\frac{\pi+2\pi}{4}\right)+i\sin\left(\frac{\pi+2\pi}{4}\right)\right)=-\sqrt{2}+i\sqrt{2}\\ \sqrt[4]n=2e^{i\frac{\pi+4\pi}{4}}=2\left(\cos\left(\frac{\pi+4\pi}{4}\right)+i\sin\left(\frac{\pi+4\pi}{4}\right)\right)=-\sqrt{2}-i\sqrt{2}\\ \sqrt[4]n=2e^{i\frac{\pi+6\pi}{4}}=2\left(\cos\left(\frac{\pi+6\pi}{4}\right)+i\sin\left(\frac{\pi+6\pi}{4}\right)\right)=\sqrt{2}-i\sqrt{2}\\ \]And those are the fourth roots of \(n\).

OpenStudy (anonymous):

Can't thank you enough! I'm on my way to make sense of it just now for myself... Thank you!

OpenStudy (anonymous):

Haha, sure thing